News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Wednesday 17 November 1999 9.00am EST

During the last 48 hours Chandra completed calibration observations of HZ-43 and the Coma cluster with HRC-S/LETG and HRC-I, and continued a GTO observation of the Hubble Deep Field with ACIS-I.

Real-time commanding was conducted overnight to perform a test with ACIS aimed at reducing the CTI degradation. The test was conducted with one chip and involved inserting pauses after each row is clocked to allow more time for trapped charge to be released. Ground testing has demonstrated improvement with this approach, and analysis will be performed on the on-orbit data to assess corresponding improvement with the flight instrument.

The team finalized preparations for the Leonid's meteor shower which is predicted to peak at 9.20pm EST 11/17. The spacecraft will be pointed directly at the anti-radiant with solar arrays clocked to minimize exposed surface area through the storm peak. In addition, a target within 4 degrees of the anti-radiant location will be observed on either side of the direct anti-radiant observation (total 21 hours). Additional DSN coverage has also been obtained and a full engineering team will be in place to monitor the real-time data overnight.

During the next 48 hours observations will be conducted of NGC 7293, the Leonid anti-radiant, MS 2137.3-2353 and PKS 2149-306.

The spacecraft subsystems continued to operate nominally.





Return to Status Reports